OpenStudy (anonymous):

Winnie rented a room at $250 per week. She also had to pay a deposit of $1200. Part A: Write a function to show the total amount of money f(n) that Winnie has to pay for renting the room for n weeks. (4 points) Part B: What is the total rent that Winnie has to pay for using the room for 12 weeks? (4 points) Part C: If the deposit amount decreases by $100, what would be the total rent for 6 weeks? (2 points)

OpenStudy (anonymous):

Part A: Total amount of money is 1200 + $250 a week for n weeks that is simply f(n) = 1200 + 250n Part B; For 12 weeks, that is 1200 + 250(12) = 1200 + 3000 = $4,200 Part C; You can now finish this on your own.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

One More? (: